# FIRE: A Failure-Adaptive Reinforcement Learning Framework for Edge Computing Migrations

### v2

@@ Abstract
In edge computing, users’ service profiles are migrated due to user mobility. Reinforcement learning (RL) frameworks have been proposed to do so, often trained on simulated data. However, existing RL frameworks overlook occasional server failures, which although rare, impact latency-sensitive applications like autonomous driving and real-time obstacle detection. Nevertheless, these failures (rare events), being not adequately represented in historical training data, pose a challenge for data-driven RL algorithms. As it is impractical to adjust failure frequency in real-world applications for training, we introduce FIRE, a framework that adapts to rare events by training a RL policy in an edge computing digital twin environment. We propose ImRE, an importance sampling-based Q-learning algorithm, which samples rare events proportionally to their impact on the value function. FIRE considers delay, migration, failure, and backup placement costs across individual and shared service profiles. We prove ImRE’s boundedness and convergence to optimality. Next, we introduce novel deep Q-learning (ImDQL) and actor critic (ImACRE) versions of our algorithm to enhance scalability. We extend our framework to accommodate users with varying risk tolerances. Through trace driven experiments, we show that FIRE reduces costs compared to vanilla RL and the greedy baseline in the event of failures.
@@ 1 Introduction
Mobile (or multi-access) edge computing (MEC) is a key technology in 5G and 6G telecommunication systems. It enables computationally intensive and latency sensitive mobile applications such as real time image processing, augmented reality, interactive gaming, etc. In MEC, computing resources such as server clusters are positioned close to end-users, typically at cellular base stations or WiFi access points at the network edge of the radio access network [1, 2]. This proximity enables latency sensitive applications, by avoiding wide-area network delays encountered in cloud computing [2].
Researchers have been actively investigating how to jointly allocate computing and radio resources to computing tasks offloaded to edge servers [1, 3, 4, 5, 6]. Nevertheless, user mobility across geographical areas presents a challenge towards task offloading [7, 8, 9, 10]. Service migration, moving service profiles to access points nearer the user as they move, has been proposed to help reduce latency and maintain service continuity [8, 9]. Yet, frequent service migration in response to user movement causes additional expenditure, due to the increase in energy consumption. Therefore, balancing the delay-migration cost tradeoff in service migration has been widely studied [7, 11, 8, 12, 13, 14, 15]. Due to the large decision space of possible migrations when the numbers of access points and time-slots increase, and the lack of information on costs and mobility patterns, markov decision processes (MDPs) and reinforcement learning (RL) are used as common solution methods [7, 11, 16, 17]. Such methods often rely on digital twins of the edge computing network to train RL models, as digital twins allow the exploration of different resource allocation policies, without generating real-world consequences.
@@ 1.1 Challenges: Edge Computing and Resilience
RL-based solutions to resource management in edge computing rely on a realistic simulation of the edge computing environment. Few migration works, however, account for another challenge in edge computing: the resilience of edge computing systems to rare but serious events like server failures. They occur due to reasons such as system overload, hardware failures, or malicious attacks, and they can be costly: the average cost of an outage at a cloud data center, for example, has increased to \$$\$740000$ in $2016$ [18]. Edge server failures may be even more probable than those in the cloud due to their distributed geographical locations, which complicates management and maintenance [19], though they will likely remain rare overall. Even worse, however, edge computing failures can have a sizeable impact as many edge computing applications are latency sensitive, such as the Internet of vehicles, augmented reality, and video processing applications [1]. Managing resources without considering rare but severe failure increases latency and reduces reliability. For instance, if the user’s service profile is migrated to an access point that then experiences a server failure, its job is not able to be completed, jeopardizing the smooth and safe functioning of applications, such as autonomous driving and real time obstacle detection.
Therefore, we need a technique to handle failures, to ensure uninterrupted application operation. We propose the use of backups, which can take over when the primary service’s edge server fails. Nevertheless, managing the migration of both the primary and backup services is challenging, especially as the number of access points and timeslots increases. Additionally, the network operator has a lack of knowledge of the user mobility distributions and associated costs. As mentioned above, MDP and RL has traditionally been proposed as a method to solve the service migration problem [7, 11, 16, 17]. Nevertheless, rare events and failures occur at a low probability, making it difficult to jointly plan or learn an optimal resource allocation policy for both the usual and rare event scenarios. This is because reinforcement learning’s reliance on past reward data may overlook the significance of rare events, impacting the training process. [20]. Therefore, we introduce FIRE: Failure-adaptive Importance sampling for Rare Events, a resilience framework for edge computing service migration.
@@ 1.2 FIRE: Failure-adaptive Importance sampling for Rare Events
Our FIRE framework employs a digital twin - inspired setup, employing reinforcement learning to optimize service migration and the backup migration in edge computing systems, amidst potential rare service failures. A digital twin edge computing system virtually replicates a physical edge computing system and mirrors its network conditions, akin to similar setups used for data center management and remote monitoring [21]. Our framework FIRE further models limited edge server capacity and communication delays to allow the exploration of service and backup migration optimization in the network, in light of rare event states. Our optimization balances the trade-offs amongst the following costs: communication and computing delay costs of the multiple users, the migration cost of service profiles, the backup storage and migration costs, along with the failure costs. Because the actual rare event probability is low and unable to help us learn an optimal policy to prepare for rare events, we use importance sampling of rare events in our digital twin setup to increase the sampling of the rare events in the reinforcement learning environment. We sample at a rate which represents their contribution to system wide costs, to help us prepare for rare events. Error correction is done through the use of importance sampling weights, to enable us to learn an optimal policy with respect to the true rare events probability. Unlike online training, training of RL in the edge computing digital twin also enables the system to learn policies while avoiding the large failure costs such as the high cost experienced by latency sensitive applications during server failures. The converged policy trained by the simulator is then applied to online scenarios, in which rare events happen at their natural probabilities.
@@
We model two decision making scenarios. Firstly, we suppose that every user has an individual service profile for job computation. Secondly, we consider multiple users sharing service profiles, for example common game environments when users are playing the same mobile or AR game together, or users sharing common neural networks for multi-modal learning [22]. The service profile replicas can also serve as backups in this second scenario [23], while in the first scenario we must create an individual backup for each user. We further consider that edge computing users may exhibit varying risk tolerances due to differences in their computing applications and latency requirements. Some users, for example, may have a higher risk tolerance for edge server failures due to their application being less latency sensitive than others, or may be less willing to incur a cost of placing and migrating backups to prepare for failures. In light of this, we propose an algorithm that caters to users of different risk tolerances, without requiring separate RL algorithm training for different risk tolerances.
In summary, the contributions of this paper are as follows:
@@
  • We present a model of service migration and backup placements given user mobility that introduces the modelling of server failures as rare events. We model the setting where all users have individual service profiles, and the setting where users share service profiles.
@@
  • We introduce FIRE, a resilience framework designed to handle rare events like server failures. As a digital twin, our importance sampling based Q-learning algorithm ImRE simulates rare events frequently without real-world consequences, allowing us to learn a failure-aware migration policy with reinforcement learning. We prove that this algorithm is bounded (Theorem 2), and that it converges to the optimal policy (Theorem 4).
@@
  • We propose deep Q-learning (ImDQL) and actor critic (ImACRE) versions of our importance sampling-based rare events adaptive algorithm, to handle large and combinatorial state and action spaces in real-world networks. These algorithms differ from traditional deep Q-Learning and actor critic as we incorporate importance sampling and error correcting weights. Additionally, we present the RiTA algorithm that adjusts service migration and backup placement based on individual risk tolerance.
@@
  • Finally, we provide trace-driven simulation results showing the convergence of our algorithms to optimality. We show that unlike vanilla Q-learning, our algorithm is able to be resilient towards server failures, resulting in sizeable cost reductions. Our importance sampling framework is able to be extended and modified, to tackle other resource allocation and decision making problems in light of rare events in networking and communications.
@@ 2 Background and Related Work
Building on the literature on service and virtual machine (VM) migration in follow-me-cloud [24, 25], service or VM placement and migration in light of user mobility in edge computing has been a widely studied problem [9, 7, 16, 11, 8, 12, 13, 14, 15]. These are often NP-hard knapsack-like problems. In particular, [7, 11] put forth an MDP formulation to make optimal migration decisions in light of unknown mobility information, with [11] using dynamic programming and reinforcement learning to find the optimal solution. [8] optimized the performance-cost tradeoff in service migration in the long run via Lyapunov optimization. [13] formulated a personalized performance service migration problem as a multi-armed bandit problem. Nevertheless, these service migration works have not approached the problem from the angle of resilience, preparing for and being adaptive to server failures.
@@
At the same time, other works have realized that cloud and edge computing presents resilience challenges, particularly regarding resource allocation [26]. Ultra-reliable offloading mechanisms have been proposed to address this challenge using two-tier optimization [27]. Proactive failure mitigation has also been proposed for edge-based NFV [28]. Failure aware workflow schedulling has been proposed for cloud computing [29]. Others have examined the use of reinforcement learning for replica placement in edge computing [30] and caching placement [31]. The migration of users at overlapping coverage areas to other servers has been considered in [32], but under this scheme, the users in the center of coverage areas still experience a higher latency as they are connected to the cloud. [18] used graphical models to learn spatio-temporal dependencies between edge server and link failures. Nevertheless, the above works did not consider resource allocation in light of unknown user mobility across different locations, or the cost of service migration. Our prior work [33] considers rare events in service migration, for a single user scenario, introducing an actor critic importance sampling algorithm. In this work, we extend the framework to consider more intricate multiple user scenarios, including individuals with both individual and shared service profiles. We also introduce tabular and deep Q-learning importance sampling algorithms, and provide theoretical guarantees for the tabular algorithm.
@@
Other problems in edge computing have been investigated from the risk-awareness perspective as well, catering to heterogeneous users and applications, for example risk-aware non-cooperative MEC offloading [34], risk-aware energy scheduling in edge computing [35], and energy budget risk-aware application placement [36]. However, they do not consider the effect of rare events on learned RL policies.
@@
(Deep) reinforcement learning has been used to optimize migration and offloading in edge computing [37, 38, 7, 11, 16, 17], with the goal of minimizing migration costs and satisfying edge resource constraints while maximizing quality-of-computation [37] or trading off the expended energy and latency [38]. Importance sampling is widely used in reinforcement learning for policy evaluation (i.e., finding the expected future reward from following a given policy $\pi$) [39], and prior work has proposed to utilize importance sampling to over-sample rare events for policy evaluation [20]. However, these two works involve estimating the value function given a fixed policy, and does not consider learning the optimal policy $\pi^{*}$ which entails policy changes. Their proof techniques do not generalize to the $Q$-learning  [40] that forms the core of FIRE. Recent work has proposed using importance sampling to enhance RL’s data efficiency [41] in more general settings, but they do not consider rare events explicitly and use a technique that requires knowledge of the full system dynamics.
3 Service Migration of Individual Service Profiles in light of Rare Events
@@
In this section, we introduce our failure-aware service migration model, where users have their individual service profiles. In the next section, we will consider users that share service profiles.
@@ 3.1 Service Placement Model
Users and edge servers. We consider an MEC system with $K$ users and $N$ access points (APs), such as a base station or wifi-access point. Each access point is equipped with a server to which users can offload their time-sensitive computing jobs, and we associate each access point with a disjoint physical region; users in this region then connect to this access point. Users are mobile and move from region to region at different times of the day. For example, office workers may move to the city districts for work during morning rush hours. This influences which access point they will be nearest to. User $k$’s location at time $t$ is represented by $l_{u}^{k}(t)$. We consider that each user will be associated with its nearest access point, via local area network (LAN). The users’ mobility pattern follows a probability distribution $m(l_{u}^{\prime k}|l_{u}^{k})$, where $m(l_{u}^{\prime k}|l_{u}^{k})$ represents the transition probability the user travels to location $l_{u}^{\prime k}$, after being at location $l_{u}^{k}$.
Each user maintains a service profile at the edge servers, e.g., in virtual machines or containers [8, 42], that consists of the state needed to run its edge computing application. Our goal is then to decide where the service profile and any backups should be located. To maintain low latencies for time-sensitive service applications, in light of user mobility, services are pre-migrated - likely to nearby APs. Service migration is stateful, involving the migration of the set of parameters associated with the user, along with the computing job’s intermediate results. At each time-slot $t$, the user sends a service request to the network operator. The location of the user’s service at time $t$ is represented by $l_{s}^{k}(t)$, and due to pre-migrations the network operator makes this placement decision in advance, before knowledge of the user location at time $t$ is known.
Failure model. In reality, rare events (system anomalies) such as server failures or shutdowns occur. Such events will impact the quality of service the user receives, as the user is not able to have his or her job computed on time, or at all, impacting the safe and smooth functioning of edge applications and jeopardizing the low latency benefits of edge computing.
@@
We formulate a Markov Decision Process (MDP) in which the system state $s(t)\in S$ is $[s_{1}(t),..s_{k}(t),...s_{K}(t)]$, where
@@
Transitions between states are governed by users’ movements from one access point ($l_{u}^{k}(t)$) to another ($l_{u}^{k}(t+1)$) following their mobility distribution $m(l_{u}^{\prime}|l_{u})$. To quantify these transitions, we let
@@
refers to the state record without the failure indicator $f_{ind}$. We let the set of these ‘state records’ be $\tilde{S}$. At each state $s$, there is a small probability of $\epsilon(s)$ (for example $\epsilon(s)=0.01$), that in the next timeslot, a server failure occurs at at least one location. The failure indicator $f_{ind}$ is $1$ when the location of at least one service profile $l_{s}^{k}(t)$ experiences a server failure. These states belong to the set of “Rare event states”, $T\subset S$, defined as follows:

### v3

@@ Abstract
In edge computing, users’ service profiles are migrated between edge servers due to user mobility. Reinforcement learning (RL) frameworks have been proposed to do so, often trained on simulated data. However, existing RL frameworks overlook occasional server failures, which although rare, impact latency-sensitive applications like AR/VR and real- time obstacle detection. These rare failures, being not adequately represented in historical training data, pose a challenge for data-driven RL algorithms. We introduce FIRE, a framework that adapts to rare events by training a RL policy in an edge computing digital twin environment. We propose FIRE-ImRE, an importance sampling-based Q-learning algorithm, which samples rare events proportionally to their impact on the value function. FIRE considers delay, migration, failure, and backup placement costs across individual and shared service profiles. We prove FIRE-ImRE’s boundedness and convergence to optimality. Next, we introduce novel deep Q-learning (FIRE-ImDQL) and actor critic (FIRE-ImACRE) versions of our algorithm to enhance scalability. We extend our framework to accommodate users with varying risk tolerances of rare failure events. Through trace-driven experiments, we show that FIRE reduces edge computing costs compared to vanilla RL and the greedy baseline in the event of failures.
@@ 1 Introduction
Mobile (or multi-access) edge computing (MEC) is a key technology in 5G and 6G telecommunication systems. It enables computationally intensive and latency sensitive mobile applications such as real time image processing, augmented reality, interactive gaming, etc. In MEC, computing resources such as server clusters are positioned close to end-users, typically at cellular base stations or WiFi access points at the edge of the radio access network [1, 2]. This proximity reduces response latency by avoiding the wide-area network delays encountered in cloud computing [2].
Researchers have been actively investigating how to jointly allocate computing and radio resources to computing tasks offloaded to edge servers [1, 3]. Nevertheless, user mobility across geographical areas presents a challenge towards task offloading [4, 5, 6]. Service migration, or moving service profiles to access points nearer the user as they move, has been proposed to maintain service continuity [5, 6]. Yet, frequent service migration causes additional expenditure, due to the required network usage. Therefore, balancing the delay-migration cost tradeoff has been widely studied [4, 7, 5, 8, 9]. Due to the large decision space when the numbers of access points and time-slots increase, and the lack of information on costs and mobility patterns, reinforcement learning (RL) is a common solution method [4, 7, 10]. Such methods often rely on digital twins of the edge computing network to train RL models, as digital twins allow the exploration of different resource allocation policies, without generating real-world consequences.
@@ 1.1 Challenges: Edge Computing and Resilience
Few migration works, however, account for another challenge in edge computing: the resilience of edge computing systems to rare but serious events like server failures. They occur due to reasons such as system overload, hardware failures, or malicious attacks, and they can be costly: the average cost of an outage at a cloud data center, for example, has increased to \$$\$740000$ in $2016$ [11]. Edge server failures may be even more probable than those in the cloud due to their distributed geographical locations, which complicates management and maintenance [12], though they will likely remain rare overall. Even worse, edge computing failures can have a sizeable impact as many edge computing applications are latency sensitive, such as the Internet of vehicles, augmented reality, and video processing applications [1]. Managing resources without considering rare but severe failure increases latency and reduces reliability. For instance, if the user’s service profile is migrated to a server that then fails, its job is not able to be completed, jeopardizing the smooth and safe functioning of applications, such as autonomous driving and real time obstacle detection.
We propose to use backups, which can take over when the primary service’s edge server fails, to handle failures. Nevertheless, managing the migration of both the primary and backup services is challenging, as the number of access points and timeslots increases. As mentioned above, due to the large decision space, and the network operator’s lack of information, RL has traditionally been proposed to solve the service migration problem [4, 7, 10, 13]. Nevertheless, rare events and failures occur at a low probability, making it difficult to jointly plan or learn an optimal resource allocation policy for both the usual and rare event scenarios. This is because reinforcement learning’s reliance on past reward data may overlook the significance of rare events, impacting the training process [14]. Therefore, we introduce FIRE: Failure-adaptive Importance sampling for Rare Events, a resilience framework for edge computing service migration.
@@ 1.2 FIRE: Failure-adaptive Importance sampling for Rare Events
Our FIRE framework employs a digital twin - inspired setup, employing RL to optimize service migration and the backup migration in edge computing systems, amidst potential rare service failures. A digital twin edge computing system virtually replicates a physical edge computing system and mirrors its network conditions, akin to similar setups used for data center management and remote monitoring [15]. Our framework FIRE further models limited edge server capacity and communication delays, and our optimization balances the trade-offs amongst the following costs: communication and computing delay costs of the multiple users, the migration cost of service profiles, the backup storage and migration costs, along with the failure costs.
Because the actual rare event probability is low and unable to help us learn an optimal policy to prepare for rare events, we use importance sampling of rare events in our digital twin setup to increase the sampling of the rare events in the RL environment. Error correction is done through the use of importance sampling weights, to enable us to learn an optimal policy with respect to the true rare events probability. Unlike online training, training of RL in the edge computing digital twin also enables the system to learn policies while avoiding the large failure costs of experiencing actual server failures. The converged policy trained by the simulator is then applied to online scenarios, in which rare events happen at their natural probabilities. While [16, 14] also use importance sampling to sample rare events, these two works involve estimating the value function given a fixed policy, and their algorithm is not applied in the edge computing scenario. In contrast, FIRE learns, not just evaluates, the optimal policy $\pi^{*}$. We further propose novel boundedness and convergence proofs. To the best of our knowledge, we are the first to propose importance sampling of rare events to prepare for failures in edge computing.
@@
We model two decision making scenarios. Firstly, we suppose that every user has an individual service profile for job computation. Secondly, we consider multiple users sharing service profiles, for example common game environments when users are playing the same mobile or AR game together, or common neural networks for multi-modal learning [17]. The service profile replicas can also serve as backups in this second scenario [18], while in the first scenario we must create an individual backup for each user. We further consider that edge computing users may exhibit varying risk tolerances due to differences in their computing applications and latency requirements, or different willingness to incur the backup placement or migration costs.
In summary, our contributions are as follows:
@@
  • We present a model of service migration and backup placements given user mobility that introduces server failures as rare events. We model both the individual and shared service profile settings.
@@
  • We introduce FIRE, a resilience framework designed to handle rare events like server failures. As a digital twin, our importance sampling based Q-learning algorithm FIRE-ImRE simulates rare events frequently without real-world consequences, allowing us to learn a failure-aware migration policy. We prove that this algorithm is bounded (Theorem 2), and that it converges to the optimal policy (Theorem 4).
@@
  • We propose deep Q-learning (FIRE-ImDQL) and actor critic (FIRE-ImACRE) versions of FIRE, to handle large and combinatorial state and action spaces in real-world networks. These algorithms differ from traditional deep Q-Learning and actor critic as we incorporate importance sampling and error correcting weights. Additionally, we present the RiTA algorithm that adjusts service migration and backup placement based on individual risk tolerance, without separate RL training for different risk tolerances.
@@
  • Finally, we provide trace-driven simulation results showing the convergence of our algorithms to optimality. We show that unlike vanilla Q-learning, FIRE’s variants are resilient towards server failures, resulting in sizeable cost reductions. Our importance sampling framework can be extended to solve other resource allocation and decision making problems in light of rare events in networking scenarios.
@@ 2 Background and Related Work
Building on the literature on service and virtual machine (VM) migration in follow-me-cloud, service or VM placement and migration in light of user mobility in edge computing has been a widely studied problem [6, 4, 10, 7, 5, 8, 9, 19]. These are often NP-hard knapsack-like problems. In particular, [4, 7] used dynamic programming and RL to make migration decisions given unknown mobility information. [5] optimized the long term performance-cost trade-off via Lyapunov optimization. Nevertheless, these service migration works have not addressed resilience, including preparation for and adaptation to server failures.
@@
Other works have realized that cloud and edge computing presents resilience challenges to resource allocation, including ultra-reliable offloading mechanisms using two-tier optimization [20], proactive failure mitigation for edge-based NFV [21], failure aware workflow scheduling for cloud computing [22], RL for replica placement in edge computing [23]. and graphical models to learn spatio-temporal dependencies between edge server and link failures [11]. The above works did not consider resilience in light of user mobility. The migration of users at overlapping coverage areas has been considered in [24], but here, users at the center of coverage areas still experience a higher latency as they are connected to the cloud. Nevertheless, the above works did not consider unknown user mobility and unknown costs. Our prior work [25] addressed rare events in single-user service migration using an actor-critic importance sampling algorithm. Here, we extend this to multi-user scenarios with both individual and shared service profiles, introduce tabular and deep Q-learning algorithms, and provide theoretical guarantees for the tabular algorithm.
@@
Other edge computing challenges have been explored from the risk-awareness perspective, catering to heterogeneous users and applications, for example risk-aware non-cooperative MEC offloading [26], and energy budget risk-aware application placement [27]. However, they do not consider the effect of rare events on learned RL policies.
@@
(Deep) RL has been used to optimize migration and offloading in edge computing [4, 26, 7, 10, 13] with the goal of optimizing across trade-offs such as migration cost, latency and energy costs. Importance sampling has been used in RL for policy evaluation (i.e., finding the expected future reward from following a given policy $\pi$) [16, 14]. However, these two works involve estimating the value function given a fixed policy, instead of learning the optimal policy $\pi^{*}$ over time. Their proof techniques do not generalize to the $Q$-learning  [28] that forms the core of FIRE. Recent work has proposed using importance sampling to enhance RL’s data efficiency [29] in general, but they do not consider rare events explicitly and require knowledge of the full system dynamics.
3 Migration of Individual Service Profiles
@@
In this section, we introduce our failure-aware service migration model, with individual service profiles. In the next section, we will consider shared service profiles.
@@ 3.1 Service Placement Model
Users and edge servers. We consider an MEC system with $K$ users and $N$ access points (APs), such as a base station or wifi-access point. Each AP is equipped with a server to which users can offload latency-sensitive computing jobs, and we associate each AP with a disjoint physical region; users in this region then connect to this access point. Users are mobile and move from region to region at different times of the day. For example, office workers may move to the city for work during morning rush hours, changing their nearest AP. User $k$’s location at time $t$ is represented by $l_{u}^{k}(t)$. We consider that each user will be associated with its nearest AP, via local area network (LAN). The users’ mobility pattern follows a probability distribution $m(l_{u}^{\prime k}|l_{u}^{k})$, where $m(l_{u}^{\prime k}|l_{u}^{k})$ represents the transition probability the user travels to location $l_{u}^{\prime k}$, after being at location $l_{u}^{k}$.
Each user maintains a service profile at the edge servers, e.g., in virtual machines or containers [5], that consists of the state needed to run its edge computing application. Our goal is then to decide where the service profile and any backups should be located. To maintain low latencies for time-sensitive service applications, in light of user mobility, services are pre-migrated - likely to nearby APs. Service migration is stateful, involving the migration of the set of parameters and intermediate results associated with the user’s computing job. At each time-slot $t$, the user sends a service request to the network operator. The location of the user’s service at time $t$ is represented by $l_{s}^{k}(t)$, and due to pre-migrations the network operator makes this placement decision before the user location at time $t$ is known.
Failure model. In reality, rare events (system anomalies) such as server failures or shutdowns occur. Such events will impact the user’s quality of service, e.g., if the user is not able to have his or her job computed on time, impacting the safe and smooth functioning of edge applications and jeopardizing the low latencies of edge computing.
@@
We formulate a Markov Decision Process (MDP) with system state $s(t)\in S$ is $[s_{1}(t),..s_{k}(t),...s_{K}(t)]$, where
@@
Transitions between states are governed by users’ movements between APs ($l_{u}^{k}(t)$ and $l_{u}^{k}(t+1)$) following their mobility pattern $m(l_{u}^{\prime}|l_{u})$. To quantify them, we let
@@
refers to the state record without the failure indicator $f_{ind}$. We let the set of these records be $\tilde{S}$. At each state $s$, there is a small probability $\epsilon(s)$ (e.g., $\epsilon(s)=0.01$), that in the next timeslot, a server failure occurs in at least one location. The failure indicator $f_{ind}$ is then set to $1$. These states belong to the set of “Rare event states”, $T\subset S$, defined as follows: